The product of two consecutive numbers is 72. Find the two integers. Can someone show me the steps to reach to the answer? Thank you.
Answer by scott8148(6628)

n (n + 1) = 72

n^2 + n - 72 = 0

(n + 9) (n - 8) = 0

n + 9 = 0 ___ n = -9

n - 8 = 0 ___ n = 8

the numbers are -9 and -8 ___ OR ___ 8 and 9